# Harmonizing Generalization and Personalization in Federated Prompt Learning

### v1

@@ Abstract
Federated Prompt Learning (FPL) incorporates large pre-trained Vision-Language models (VLM) into federated learning through prompt tuning. The transferable representations and remarkable generalization capacity of VLM make them highly compatible with the integration of federated learning. Addressing data heterogeneity in federated learning requires personalization, but excessive focus on it across clients could compromise the model’s ability to generalize effectively. To preserve the impressive generalization capability of VLM, it is crucial to strike a balance between personalization and generalization in FPL. To tackle this challenge, we proposed Federated Prompt Learning with CLIP Generalization and low-rank Personalization (FedPGP), which employs pre-trained CLIP to provide knowledge-guidance on the global prompt for improved generalization and incorporates a low-rank adaptation term to personalize the global prompt. Further, FedPGP integrates a prompt-wise contrastive loss to achieve knowledge guidance and personalized adaptation simultaneously, enabling a harmonious balance between personalization and generalization in FPL. We conduct extensive experiments on various datasets to explore base-to-novel generalization in both category-level and domain-level scenarios with heterogeneous data, showing the superiority of FedPGP in balancing generalization and personalization.
@@ 1 Introduction
  • We are the first to consider both personalization and generalization in federated prompt learning. We aim to learn a personalized prompt for each client in heterogeneous federated scenarios while preserving the remarkble generalization capacity in Visual-Language models, leading to the balance of generalization and personalization.
@@ 2 Related Work
Federated Prompt Learning Prompt tuning is a technique employed to adapt pre-trained models to diverse downstream tasks. For instance, CoOp (Zhou et al., 2022b) uses tunable text prompts to replace the fixed template in CLIP, and CoCoOp (Zhou et al., 2022a) utilizes image feature to instruct the optimization of the soft text prompt. ProGrad (Zhu et al., 2023) selectively updates prompts based on aligned gradients with general knowledge to prevent forgetting essential information from VLMs. Some works also incorporate prompt tuning into federated learning. PromptFL (Guo et al., 2023b) introduced prompt learning into Federated Learning. FedPR (Feng et al., 2023) focuses on learning federated visual prompts within the null space of the global prompt for MRI reconstruction. pFedPG (Yang et al., 2023) employs a client-specific prompt generator on the server side for personalized promptsm, while FedTPG (Qiu et al., 2023) also trains a global prompt generation network to enhance generalization. pFedprompt (Guo et al., 2023a) maintains a non-parametric personalized attention module for each client to generate local personalized spatial visual features. FedOTP (Li et al., 2024) employs unbalanced Optimal Transport to promote the collaboration between global and local prompts across heterogeneous clients. Designed for domain discrepancy, FedAPT (Wei et al., 2023) unlocks specific domain knowledge for each test sample to provide personalized prompts, and Fed-DPT (Su et al., 2022) applies both visual and textual prompt tuning to facilitate domain adaptation over decentralized data. However, these methods overlook the aspect of generalization. We propose FedPGP, a framework that effectively balances personalization and generalization in FPL.
@@
Contrastive learning Contrastive learning methodologies have gained significant attention by consistently attaining state-of-the-art outcomes results in the field of visual representation learning (Chen et al., 2020a, b; Xie et al., 2022). The fundamental principle behind contrastive learning is minimize the distance between representations generated from diverse augmentations of the same image (positive pairs) while simultaneously maximizing the distance between representations obtained from augmented views of different images (negative pairs). A proportion of research (Tan et al., 2022b; Li et al., 2021a; Mu et al., 2023) combines contrastive learning with federated learning, which improves the local training process and achieves higher model effectiveness. FedPCL (Tan et al., 2022b) employs prototype-wise contrastive learning for client-specific representations, promoting alignment with global and local prototypes to enhance knowledge sharing. MOON (Li et al., 2021a) minimizes the distance between local and global model representations while increasing the distance from the previous local model’s representation. Different from model-wise of MOON and prototype-wise of FedPCL, our FedPGP introduces a novel prompt-wise contrastive methodology. In addition, unlike previous work focusing on aligning the global and local components, FedPGP treats representations of global and personalized prompts as negative pairs and representations of global and handcrafted prompts as positive pairs.
@@ 3 Proposed Method
In this section, we delve into the details of our proposed FedPGP, illustrated in Figure 1. FedPGP leverages CLIP knowledge-guidance and low-rank adaptation with an additional contastive loss to balance generalization and personalization.
Figure 1: Pipeline of FedPGP. On the left, clients send global prompts to the server for aggregation while retaining adaptation term locally. The right shows the workflow of CLIP knowledge-guidance and low-rank adaptation with an additional contastive loss to balance generalization and personalization.
@@ 4.3 Ablation Study
Effect of Parameter $\mu$ of Contrastive Loss In this subsection, we investigated the impact of the contrastive loss parameter $\mu$ in a Pathlogical Non-IID setting across four datasets with varying shot numbers. The results are presented in Figure 2, where shows an improvement in test accuracy with an increase in the number of shots. Upon observation, optimal results are mostly achieved with $\mu$ set to 1 in experiments, leading to our adoption of $\mu=1$ for other experiments.
@@ 5 Conclusion
In this paper, we propose a novel approach named FedPGP, which represents a pioneering effort to harmonize personalization and generalization in federated prompt learning. In our approach, each client gains generalization capabilities through knowledge-guidance of CLIP and acquires personalization abilities by adapting the global prompt to a personalized prompt. Further, with low-rank decomposition adaptation and an extra contrastive loss, FedPGP learns a personalized prompt for each client in heterogeneous federated scenarios while preserving the remarkble generalization capacity in pre-trained Vision-Language models. Extensive experiments on various datasets explored base-to-novel generalization in both unseen categories and domains, showing the superiority of FedPGP in balancing generalization and personalization. In future work, we aim to explore the theoretical foundations of low-rank adaptation in federated prompt learning.
6 Acknowledgement

### v2

@@ Abstract
Federated Prompt Learning (FPL) incorporates large pre-trained Vision-Language models (VLM) into federated learning through prompt tuning. The transferable representations and remarkable generalization capacity of VLM make them highly compatible with the integration of federated learning. Addressing data heterogeneity in federated learning requires personalization, but excessive focus on it across clients could compromise the model’s ability to generalize effectively. To preserve the impressive generalization capability of VLM, it is crucial to strike a balance between personalization and generalization in FPL. To tackle this challenge, we proposed Federated Prompt Learning with CLIP Generalization and low-rank Personalization (FedPGP), which employs pre-trained CLIP to provide knowledge-guidance on the global prompt for improved generalization and incorporates a low-rank adaptation term to personalize the global prompt. Further, FedPGP integrates a prompt-wise contrastive loss to achieve knowledge guidance and personalized adaptation simultaneously, enabling a harmonious balance between personalization and generalization in FPL. We conduct extensive experiments on various datasets to explore base-to-novel generalization in both category-level and domain-level scenarios with heterogeneous data, showing the superiority of FedPGP in balancing generalization and personalization. Code is available at https://github.com/TianyuCuiOvO/FedPGP.
@@ 1 Introduction
  • We are the first to consider both personalization and generalization in federated prompt learning. We aim to learn a personalized prompt for each client in heterogeneous federated scenarios while preserving the remarkable generalization capacity in Visual-Language models, leading to the balance of generalization and personalization.
@@ 2 Related Work
Federated Prompt Learning Prompt tuning is a technique employed to adapt pre-trained models to diverse downstream tasks. For instance, CoOp (Zhou et al., 2022b) uses tunable text prompts to replace the fixed template in CLIP, and CoCoOp (Zhou et al., 2022a) utilizes image feature to instruct the optimization of the soft text prompt. ProGrad (Zhu et al., 2023) selectively updates prompts based on aligned gradients with general knowledge to prevent forgetting essential information from VLMs. Some works also incorporate prompt tuning into federated learning. PromptFL (Guo et al., 2023b) introduced prompt learning into Federated Learning. FedPR (Feng et al., 2023) focuses on learning federated visual prompts within the null space of the global prompt for MRI reconstruction. pFedPG (Yang et al., 2023) employs a client-specific prompt generator on the server side for personalized prompts, while FedTPG (Qiu et al., 2023) also trains a global prompt generation network to enhance generalization. pFedprompt (Guo et al., 2023a) maintains a non-parametric personalized attention module for each client to generate local personalized spatial visual features. FedOTP (Li et al., 2024) employs unbalanced Optimal Transport to promote the collaboration between global and local prompts across heterogeneous clients. Designed for domain discrepancy, FedAPT (Wei et al., 2023) unlocks specific domain knowledge for each test sample to provide personalized prompts, and Fed-DPT (Su et al., 2022) applies both visual and textual prompt tuning to facilitate domain adaptation over decentralized data. However, these methods overlook the aspect of generalization. We propose FedPGP, a framework that effectively balances personalization and generalization in FPL.
@@
Contrastive learning Contrastive learning methodologies have gained significant attention by consistently attaining state-of-the-art outcomes results in the field of visual representation learning (Chen et al., 2020a, b; Xie et al., 2022). The fundamental principle behind contrastive learning is to minimize the distance between representations generated from diverse augmentations of the same image (positive pairs) while simultaneously maximizing the distance between representations obtained from augmented views of different images (negative pairs). A proportion of research (Tan et al., 2022b; Li et al., 2021a; Mu et al., 2023) combines contrastive learning with federated learning, which improves the local training process and achieves higher model effectiveness. FedPCL (Tan et al., 2022b) employs prototype-wise contrastive learning for client-specific representations, promoting alignment with global and local prototypes to enhance knowledge sharing. MOON (Li et al., 2021a) minimizes the distance between local and global model representations while increasing the distance from the previous local model’s representation. Different from model-wise of MOON and prototype-wise of FedPCL, our FedPGP introduces a novel prompt-wise contrastive methodology. In addition, unlike previous work focusing on aligning the global and local components, FedPGP treats representations of global and personalized prompts as negative pairs and representations of global and handcrafted prompts as positive pairs.
@@ 3 Proposed Method
In this section, we delve into the details of our proposed FedPGP, illustrated in Figure 1. FedPGP leverages CLIP knowledge-guidance and low-rank adaptation with an additional contrastive loss to balance generalization and personalization.
Figure 1: Pipeline of FedPGP. On the left, clients send global prompts to the server for aggregation while retaining adaptation term locally. The right shows the workflow of CLIP knowledge-guidance and low-rank adaptation with an additional contrastive loss to balance generalization and personalization.
@@ 4.3 Ablation Study
Effect of Parameter $\mu$ of Contrastive Loss In this subsection, we investigated the impact of the contrastive loss parameter $\mu$ in a Pathological Non-IID setting across four datasets with varying shot numbers. The results are presented in Figure 2, which shows an improvement in test accuracy with an increase in the number of shots. Upon observation, optimal results are mostly achieved with $\mu$ set to 1 in experiments, leading to our adoption of $\mu=1$ for other experiments.
@@ 5 Conclusion
In this paper, we propose a novel approach named FedPGP, which represents a pioneering effort to harmonize personalization and generalization in federated prompt learning. In our approach, each client gains generalization capabilities through knowledge-guidance of CLIP and acquires personalization abilities by adapting the global prompt to a personalized prompt. Further, with low-rank decomposition adaptation and an extra contrastive loss, FedPGP learns a personalized prompt for each client in heterogeneous federated scenarios while preserving the remarkable generalization capacity in pre-trained Vision-Language models. Extensive experiments on various datasets explored base-to-novel generalization in both unseen categories and domains, showing the superiority of FedPGP in balancing generalization and personalization. In future work, we aim to explore the theoretical foundations of low-rank adaptation in federated prompt learning.
Acknowledgement